OpenStudy (anonymous):

Use foil method to find the product (6x^5-9)(x^7-9)

OpenStudy (amistre64):

ugh.... no foil

OpenStudy (amistre64):

6x^5 -9 x^7 -9 --------- 6x^12 -9x^7 -54x^5 +81 ---------------------- \(6x^{12}-9x^7-54x^5+81\)

OpenStudy (anonymous):

\[(6x^5-9)(x^7-9)\] \[=6x^5\times x^7 + 6x^5\times -9 + (-9)\times x^7 + (-9)\times (-9)\]

OpenStudy (anonymous):

= what amistre said

OpenStudy (amistre64):

:) and i dint even have to take the wrapper off

OpenStudy (anonymous):

94 and counting. but for some reason people write "thank you" but then do not click on "good answer" and so the problems stay current. need to get with the program

OpenStudy (amistre64):

theres a good answer button?

OpenStudy (anonymous):

aswarr i had a typo.

OpenStudy (anonymous):

I do say thank you for the all the time because that is southern thing to do and yes most of the time I do click good answer

OpenStudy (anonymous):

oh i was not referring to you!

OpenStudy (anonymous):

i was conferring with my learned colleague amistre

OpenStudy (amistre64):

down here in florida we just stare blankly at each other and grin lol

OpenStudy (anonymous):

up here in the north we just flip each other off and keep walking

OpenStudy (amistre64):

banjo music plays inthe background..

OpenStudy (anonymous):

no squealing!

OpenStudy (anonymous):

well Im in tennessee originally from alabama, most of the time we speak. If not u know we are mad at something

OpenStudy (amistre64):

:)